OpenStudy (anonymous):

In how many ways can 3 singers be selected from 5 who came to an audition? A. 1 B. 10 C. 5 D. 60

OpenStudy (anonymous):

We have 5 choices for the first singer, then 4 for the 2nd and 3 for the third. So we have 5x4x3=60 choices.

OpenStudy (raden):

it is a combination problem. use 5C3 = 5!/(5-3)!3! = 5!/2!3! = 5 * 4 * 3!/2*1*3! = 10

OpenStudy (anonymous):

Oops so it is. Sorry for that!

OpenStudy (anonymous):

I need to divide by 3! since the order is irrelevant...